a) A hydraulic press has a 30mm ratio input piston and 75mm radio output piston. What input force is needed to provide an output force that is able to lift a truck 1300kg?

b) The inlet tube that supplies air pressure to operate a hydraulic jack has a 30 mm radius, The output position has a 64mm radius. What air pressure should be used to lift with the crane a 2300kg mass?


1
Expert's answer
2021-09-15T11:38:49-0400

a)


"F_{i}=mg\frac{d^2_i}{d^2_o}\\
F_{i}=(1300)(9.8)\frac{30^2}{75^2}=2040\ N"

b)


"p=\frac{2300\cdot9.8}{\pi(0.064)^2}=1.75\cdot10^6\ Pa"